OpenStudy (anonymous):

A bullet of mass 20 g travelling horizontally at 100m s- 1 embeds itself in the centre of a block of wood of mass 1 kg which is suspended by light vertical strings 1 m in length. Calculate the maximum inclination of the strings to the vertical.

OpenStudy (amistre64):

potential energy might prove useful

OpenStudy (amistre64):

the total energy in the system at the start never changes, and assuming a perfect world, all the energy is stored into the block of wood at its highest point

OpenStudy (amistre64):

1/2 mv^2 = mgh 1/2 v^2 = gh 1/2g v^2 = h 

OpenStudy (amistre64):

cos(a) = 1-h a = arccos (1-h)

OpenStudy (amistre64):

i do not provide solutions, i provide a process that will get you to the solution. I have defined all that needs to be defined, and the information for the specifics is already presented to you in the problem. it is up to you to fill in the results, or ask questions about what you do not understand instead of expecting me or anyone else to do all the work for you.